OpenStudy (anonymous):

A circular disc has a diameter of 24 centimeters. What is the area of the disc?

OpenStudy (anonymous):

144pi

OpenStudy (rulnick):

144 pi cm^2, because area = pi d^2 / 4 and d = 24 cm.

OpenStudy (anonymous):

or... simply A=pi(r^2) and r=d/2

OpenStudy (rulnick):

In decimal, that's about 452.4 cm^2.

OpenStudy (chriss):

\[\pi*r^{2}\] \[\pi*12^{2}\] \[144\pi\]